Grade 11Mechanics

The distance d of a particle is proportional to the square of the time t .The acceleration of the particle is proportional to ?

Neelabh Jaiswal
8 Years agoGrade 11

1 Answer

Harsh jain
8 Years ago
S is propotional to t ^2S = k t^2a = ??a = d^2s/dt^2a=k Since a is independent of timeAs acceleration is double differentiation of distance